# An adaptive finite element multigrid solver using GPU acceleration

### v2

@@ 6 Conclusions
We are planning several steps for further optimisation: A major limitation of the previous GPU parallelization is that the matrices are assembled on the CPU and then transferred statically to the GPU. This severely restricts the applicability to nonlinear problems. The obvious option is to assemble the matrices directly on the GPU. However, this is complicated by a large number of options in the choice and control of the discretization and the use of adaptive grids. Alternatively, we plan to use the GPU and CPU in hybrid mode, so that for nonlinear problems the Jacobi matrix of the Newton solver is always built in the background on the CPU and then transferred to the GPU. This can be seamlessly integrated in the usual inexact Newton algorithm, which only reassembles the Jacobian when the convergence rate deteriorates. A further important step is the GPU parallelization of more powerful multigrid smoothers than the block Jacobi iteration. The realisation of a Vanka smoother can be easily implemented with custom kernels and then allows the direct solution of saddle point problems.
@@
The source code of the Gascoigne library with CUDA support is found in the Zenodo repository https://zenodo.org/records/13891228 [26]. This repository includes all scripts required to reproduce the examples discussed in this manuscript.
@@ Appendix A Details on the implementation
nonlin_device$<$blocks, tpb$>$(n, dest, src);

### v3

@@ 6 Conclusions
We are planning several steps for further optimisation: A major limitation of the current GPU parallelization is that the matrices are assembled on the CPU and then transferred statically to the GPU. This severely restricts the applicability to nonlinear problems. The obvious option is to assemble the matrices directly on the GPU. However, this is complicated by a large number of options in the choice and control of the discretization and the use of adaptive grids. Alternatively, we plan to use the GPU and CPU in hybrid mode, so that for nonlinear problems the Jacobi matrix of the Newton solver is always built in the background on the CPU and then transferred to the GPU. This can be seamlessly integrated in the usual inexact Newton algorithm, which only reassembles the Jacobian when the convergence rate deteriorates. A further important step is the GPU parallelization of more powerful multigrid smoothers than the block Jacobi iteration. The realisation of a Vanka smoother can be easily implemented with custom kernels and then allows the direct solution of saddle point problems.
@@
The source code of the Gascoigne library with CUDA support is found in the Zenodo repository https://zenodo.org/records/13891228 [ZenodoCode]. This repository includes all scripts required to reproduce the examples discussed in this manuscript.
@@ Appendix A Details on the implementation
nonlin_device<<<blocks, tpb>>>(n, dest, src);